OpenStudy (anonymous):

1. If the height h of a triangle is 3 inches less than the length of the base b, and the area A of the triangle is 19 times the length of the base, find b and h.

OpenStudy (nory):

Base = b. Here's a picture of the triangle: 

OpenStudy (nory):

Now, area = 1/2 * base * height

OpenStudy (nory):

We don't know the base, but we know the height and the area in terms of the base. Does that make sense so far?

OpenStudy (anonymous):

mhmm...

Directrix (directrix):

h = b - 3 The area of a triangle is (1/2)bh. This triangle has area 19b. So, 19b = (1/2) * b * (b - 3) Crank out the algebra. Post what you get. I think we will then need to solve a quadratic equation to solve for b.

OpenStudy (wolf1728):

length of height (h) = base (b) -3 Area of triangle = 19 * b Area of ANY triangle = .5 * h * b

OpenStudy (wolf1728):

hmmmm seems Directrix was a bit quicker than I was

OpenStudy (anonymous):

I got \[19b= \frac{ 1 }{ 2 }*b^{2}-3b\]

OpenStudy (wolf1728):

19b = .5 * b * (b - 3) 19b = .5b² -3b seems jellyfishe has this too

OpenStudy (wolf1728):

19b = .5b² -3b .5b² -22b = 0 b= --22 +- sqrt(484 - 4 * .5 * 0)/2 * .5 b= 22 +- sqrt(484) /1 b= 22 + 22 = 44

OpenStudy (wolf1728):

so h = 41

OpenStudy (anonymous):

I have the answers ( h=38 b=41), but I can't figure out how to get them...

OpenStudy (wolf1728):

Darn - Are you sure those are the answers jellyfishe ?

OpenStudy (anonymous):

my textbook has the answers for that #

OpenStudy (wolf1728):

sheesh I'll recheck my work then

OpenStudy (anonymous):

I think I got it...

OpenStudy (wolf1728):

19b = .5b² -3b Okay I messed up here I should have divided 3b by 2 so: 19b = .5b² -1.5b .5b² -20.5b = 0 b = 41 and h = 38

OpenStudy (anonymous):

\[19b=\frac{ 1 }{ 2 }b*(b-3)\] If you divide the 19b by .5b (of course you have to do this to both side of the equal sign) then you get 38 because the b's cancel each other out... this make the equation: \[38=b-3\] You add three to both sides and get 41=b. :)

OpenStudy (wolf1728):

Yes I discovered that error too it should have been -1.5b instead of -3b

OpenStudy (anonymous):

I miss Algebra 2...

OpenStudy (wolf1728):

Then what is the course you are taking?

OpenStudy (anonymous):

Geometry

OpenStudy (wolf1728):

I like geometry - didn't like Algebra 2 that much (thought it was boring)

OpenStudy (anonymous):

I'm totally opposite. The only thing I didn't like was graphing.

OpenStudy (wolf1728):

To me, what sums up Algebra 2 was students went to the blackboard to show some answers. One kid had prblem #10 - a three unknown problem that was really involved. Truck X arrives and drops off 1/2 its load - truck Y drops off 44 pounds more than truck X dropped off...... I remember the kid put the answer for problem 10 as 10 then X ten (written) 5+5 I think that summed up EVERYONE'S feelings

OpenStudy (anonymous):

Our teacher had us do a lot of hand outs so we didn't really have to do that lol

OpenStudy (wolf1728):

that incident happened DECADES ago and I still find it hilarious

OpenStudy (anonymous):

i would too ha!

OpenStudy (wolf1728):

I was laughing as I typed it LOL

OpenStudy (wolf1728):

gotta go - nice chatting with you :-)

OpenStudy (anonymous):

same. Thanks for the help!! ;)